# RFWave: Multi-band Rectified Flow for Audio Waveform Reconstruction

### v2

@@ 3.2 Loss functions
In our preliminary experiments, we observe the presence of low-volume noise in regions that are expected to be silent. We attribute this to the property of Mean Square Error (MSE) used in (2). The MSE measures the absolute distortion between the predicted values and the ground truth. Since the values in silent regions are close to zero, even a minor absolute distortion in predictions can lead to a significant relative error. Consequently, models trained with the MSE produce small absolute distortions in silent regions, which are then perceived as noise.
@@
Within the multi-band structure, each subband is predicted independently, potentially resulting in inconsistencies among them. To mitigate these inconsistencies, we introduce an overlap loss. This involves maintaining overlaps between the subbands when dividing the full-band complex spectrograms. A detailed illustration of this scheme is provided in Figure 2. The complex spectrogram is circularly padded in the feature dimension with a size of $(d_{ol},d_{ol}-1)$, where $d_{ol}$ signifies the overlap size. The dimension of each subband’s main section, denoted as $d_{m}$, is calculated by dividing $d-1$ by the total number of subbands. Here, $d$ represents the dimension of the complex spectrograms. The last subband is an exception, having an extra feature dimension but one less padding dimension. Subbands are extracted by applying a sliding window along the feature dimension, where the window has a size of $d_{m}+2d_{ol}$ and shifts by $d_{m}$ [^3] The feature dimension of each subband, represented as $d_{s}$, equals $2(d_{m}+2d_{ol})$, accounting for the interleave of real and imaginary parts.. In this paper, we employ an 8-dimensional overlap. During the training phase, the MSEs of the overlapped predictions is minimized. In the inference phase, the overlaps are removed, and all subbands are merged to recreate the full-band complex spectrograms. As each subband is predicted, the model internally maintains consistency between its overlap section and the main section. The overlap serves as an anchor to maintain consistency among the subbands.
Figure 2: An illustration of dividing complex spectrograms into subbands. The area highlighted in purple represents a subband, while the section enclosed by the two dashed vertical lines indicates the main section.
@@
The magnitude spectrogram derived loss [28] is extensively utilized in GAN-based vocoders, such as HiFi-GAN [9] and Vocos [12], both of which leverage a Mel-spectrogram loss. Nevertheless, its use in diffusion-based vocoders is relatively rare. This might stem from its lack of direct compatibility with the formalization of a noise prediction diffusion model. Here we adopt STFT loss for RFWave. According to (5), the model’s output serves as an approximation for velocity:
@@
The STFT loss can be applied on the approximation $\overset{\sim}{X_{1}}$, incorporating the spectral convergence loss and log-scale STFT-magnitude loss as proposed in [28]. According to our experimental findings, the STFT loss effectively reduces artifacts in the presence of background noise.
4 Experiments
@@
We evaluate our model across various auditory contexts, including singing (vocal), music (sound effects), and extensive speech datasets, validating its comprehensive applicability and efficiency in reconstructing audio waveforms from both Mel-spectrograms and discrete EnCodec tokens. Then, we conduct ablation studies to assess the impact of various design choices.
@@
For waveform reconstruction from Mel-spectrograms, we use the Opencpop dataset [29] of 100 Mandarin songs for singing data, the MTG-Jamendo dataset [30] with over 55,000 tagged audio tracks for music, and the LibriTTS corpus[31], a multi-speaker English dataset with around 600 hours of diverse recordings, for speech data. When reconstructing waveforms from EnCodec tokens, we employ Common Voice 7.0 [32] and clean data from DNS Challenge 4 [33] for speech; the MTG-Jamendo dataset [30] for music; and FSD50K [34] and AudioSet [35] for environmental sounds.
For testing, we use 20 segments from Opencpop, a random selection of 300 sentences from the test-clean subset of LibriTTS, and 50 randomly selected audio files from MTG-Jamendo. When evaluating the quality of waveform reconstruction from Encodec tokens, we opt for an external dataset rather than a portion of the original. This more comprehensive and systematic approach covers as many domains and scenarios as possible. We have selected 900 test audio samples from 15 external datasets, spanning three major scenarios: speech, vocals, and sound effects. Details are provided in Table A.4.
@@
We utilize the time-domain model with the 3 enhanced loss functions as the default configuration. The RFWave backbone contains 8 ConvNeXtV2 blocks, and the complex spectrogram is divided into 8 equally spanned subbands. For evaluation purposes, we use 10 sampling steps. Table A.3 lists the parameters used for extracting Mel-spectrograms and complex spectrograms from datasets with different sample rates. Additional details regarding the implementation can be found in the Appendix Section A.1. Information concerning the required computational resources is detailed in the Appendix Section A.2.
Baseline and Evaluation Metrics
We conducted a comprehensive benchmark comparison of RFWave with several state-of-the-art methods, including Vocos [12], PriorGrad [15], and FreGrad [16], for reconstructing waveforms from Mel-spectrograms. Additionally, we compared RFWave with Multi-Band Diffusion (MBD) and the original EnCodec decoder for waveform reconstruction from Encodec tokens. For training Vocos, PriorGrad, and FreGrad, We adhered to the original training details provided by their open-sourced codes. The training was conducted separately on the LibriTTS, Opencpop, and MTG-Jamendo datasets, with the exception that a publicly available pre-trained Vocos model was utilized for the LibriTTS dataset. We also utilized publicly available pre-trained versions of MBD and EnCodec for comparison. We adhere to the sampling steps recommended by the authors for PriorGrad (6 steps), FreGrad (50 steps) and MBD (20 steps). The utilized repositories are detailed in Table A.2.
For the objective evaluation, we use ViSQOL [36] to assess perceptual quality, employing speech-mode for 22.05/24 kHz waveforms and audio-mode for 44.1 kHz waveforms. Additional metrics are also incorporated into our evaluation. Among these are the Perceptual Evaluation of Speech Quality (PESQ) [37], the F1 score for voiced/unvoiced classification (V/UV F1), and the periodicity error (Periodicity) [38]. As for the subjective evaluation, we conduct crowd-sourced assessments, employing a 5-point Mean Opinion Score (MOS) to determine the naturalness of the audio recordings. Participants are tasked with rating the speech samples on a scale from 1, denoting ’poor or completely unnatural’, to 5, representing ’excellent or completely natural’.
5 Results
Mel-specotrograms
The average MOS and objective metrics for various auditory contexts are listed in Table 1, and the separate results for each class are listed in Table A.5. RFWave outperforms the other models in most metrics. It achieves the highest scores in MOS, PESQ, and V/VU F1, and the lowest in periodicity error. Vocos achieves the highest ViSQOL, which might be misleading. In the MTG-Jamendo dataset, despite having the highest ViSQOL, Vocos exhibits a metallic artifacts that could affect the listening experience. This issue is also observed in [21] concerning GAN-based methods. The reason why RFWave obtains the best MOS may be because RFWave produces clearer and more consistent harmonics, especially in high-frequency. A collection of spectrograms and their corresponding analyses can be found in Appendix Section A.3.
@@ 5 Results
Table 1: Average Mean Opinion Score (MOS) and objective evaluation metrics for RFWave, Vocos, PriorGrad and FreGrad across various test sets. MOS is provided with 95% confidence interval
@@
Discrete EnCodec Tokens
We use Classifier-Free Guidance (CFG) [39] with a coefficient of 2.0 for waveform reconstruction from EnCodec tokens, enhancing the listening experience significantly. However, its effect is less pronounced with waveform reconstruction from Mel-spectrograms, suggesting the ample, deterministic information in Mel-spectrograms reduces CFG’s impact on the listening experience.
Table 2 displays the average MOS and objective metrics for EnCodec tokens with varying bandwidths across different auditory contexts. Meanwhile, Table A.6 offers the individual results for each category. RFWave excels in all metrics, achieving optimal scores, except for ViSQOL. While using a larger bandwidth generally improves performance, for RFWave, an increase in bandwidth from 6.0 kbps to 12.0 kbps results in only a slight enhancement in MOS. Again, it is the EnCodec’s GAN-based decoder that attains the highest ViSQOL. This may indicate that different model families produce waveforms with different footprints. ViSQL appears to have a subtle bias towards the ones produced by the GAN-based model. MBD outperforms the original EnCodec decoder in subjective evaluations, a result that aligns with the findings reported in [21]. In Table A.6, it is evident that MBD performs particularly well in the sound effect dataset. This could be attributed to it having more parameters and being specifically designed for music generation.
@@
We perform ablation studies on the LJSpeech dataset [40], with 250 sentences reserved as a test set, specifically for waveform reconstruction from Mel-spectrograms. The performance metrics of the model, both in the frequency and time domains, with and without the implementation of the energy-balanced loss, are summarized in Table 3. Despite the metrics not consistently correlating with human evaluations, they accurately captured the quality improvements brought about by design modifications [21]. The model, operating in the time domain with the energy-balanced loss, demonstrates superior performance in terms of ViSQOL, V/UV F1, and periodicity error, outperforming the other three configurations. The overlap loss, STFT loss, and CFG are tested on this configuration. STFT loss improves PESQ scores but may negatively impact ViSQOL and periodicity error. This is due to the model’s tendency to prioritize magnitude over phase when STFT loss is applied. This trade-off is necessary to effectively remove artifacts, particularly when background noise is present. CFG consistently increases performance, however, it doubles the inference computation and increases latency by approximately 50%.
Table 3: Objective metrics assess the model’s performance in frequency (freq) and time (time) domains, with and without the inclusion of energy-balanced loss (energy-balanced), classifier-free guidance with a guidance coefficient of 2.0 (CFG2), overlap loss (overlap), and STFT loss (STFT).
@@
freq w/o energy-balanced 3.872 4.430 0.948 0.144
freq w/ energy-balanced 3.901 4.433 0.950 0.140
time w/o energy-balanced 4.127 4.551 0.957 0.124
time w/ energy-balanced 4.181 4.598 0.965 0.110
    w/ overlap 4.158 4.599 0.959 0.122
    w/ overlap, STFT 4.211 4.578 0.961 0.119
    w/ overlap, CFG2 4.225 4.671 0.973 0.078
    w/ overlap, CFG2, STFT 4.240 4.663 0.972 0.081
Vocos(ISTFT) 3.787 4.675 0.961 0.108
PriorGrad 3.810 4.216 0.945 0.136
FreGrad 3.767 4.356 0.946 0.136
Inference Speed
We performe inference speed benchmark tests using an NVIDIA GeForce RTX 4090 GPU. The implementation was done in PyTorch [41], and no specific hardware optimizations were applied. The inference was carried out with a batch size of 1 sample, utilizing the LJSpeech test set. Table 4 displays the synthesis speed and model size of the models. RFWave achieves the highest inference speed among the diffusion models, making latency no longer an obstacle for application. Vocos stands as a strong baseline given that it requires only a single forward pass and operates at the frame level.

### v3

@@ 3.2 Loss functions
In preliminary experiments, we noticed low-volume noise in expected silent regions. We attribute this to the property of Mean Square Error (MSE) used in (2). The MSE measures the absolute distortion between the predicted values and the ground truth. Since the values in silent regions are close to zero, even a minor absolute distortion in predictions can lead to a significant relative error. Consequently, models trained with the MSE produce small absolute distortions in silent regions, which are then perceived as noise.
@@
Energy-balanced loss is related to PriorGrad (gil Lee et al., 2022), yet they are fundamentally different. PriorGrad uses a data-dependent diffusion prior to narrow the gap between the prior and data distribution, thereby reducing sampling steps. This prior is manually chosen based on specific datasets and tasks. In contrast, our method employs standard Gaussian noise as the diffusion prior and adjusts the loss function according to the target’s standard deviation. Importantly, our approach requires no manual selection.
@@
Within the multi-band structure, each subband is predicted independently, potentially resulting in inconsistencies among them. To mitigate these inconsistencies, we introduce an overlap loss. This involves maintaining overlaps between the subbands when dividing the full-band complex spectrograms. A detailed illustration of this scheme is provided in Figure 2 and described in detail in Appendix Section A.3. In this paper, we employ an 8-dimensional overlap. During the training phase, the MSEs of the overlapped predictions is minimized. In the inference phase, the overlaps are removed, and all subbands are merged to recreate the full-band complex spectrograms.
As each subband is predicted, the model internally maintains consistency between its overlap section and the main section. The overlap serves as an anchor to maintain consistency among the subbands. Modeling higher bands based on lower bands rather than predicting all subbands in parallel can increase consistency between subbands. However, at the inference stage, if the lower bands are incorrectly predicted, the higher bands conditioned on them would also be inaccurate (Roman et al., 2023). RFWave addresses this by using overlap loss to maintain consistency between subbands during training. At the inference stage, subbands are predicted independently, so any error occurring in one subband does not negatively affect the others.
Figure 2: An illustration of dividing complex spectrograms into subbands. The area highlighted in pink represents a subband, while the section enclosed by the two dashed vertical lines indicates the main section.
@@
The magnitude spectrogram derived loss (Arik et al., 2018) is extensively utilized in GAN-based vocoders, such as HiFi-GAN (Kong et al., 2020a) and Vocos (Siuzdak, 2023), both of which leverage a Mel-spectrogram loss. Nevertheless, its use in diffusion-based vocoders is relatively rare. This might stem from its lack of direct compatibility with the formalization of a noise prediction diffusion model. Here we adopt STFT loss for RFWave. According to (5), the model’s output serves as an approximation for velocity:
@@
The STFT loss can be applied on the approximation $\overset{\sim}{X_{1}}$, incorporating the spectral convergence loss and log-scale STFT-magnitude loss as proposed in (Arik et al., 2018), which are detailed in Appendix Section A.4. According to our experimental findings, the STFT loss effectively reduces artifacts in the presence of background noise.
3.3 Selecting Time Points for Euler Method
Liu et al. (2023) employ equal step intervals for the Euler method, as illustrated in (4) . Here, we propose selecting time points for sampling based on the straightness of transport trajectories. The time points are chosen such that the increase in straightness is equal across each step. The straightness of a learned velocity filed $v$ is defined as $S(v)=\int_{0}^{1}E\mid\mid(X_{1}-X_{0})-v(X_{t},t\mid\mathcal{C})\mid\mid^{2}\mathrm{d}t$, which describes the deviation of the velocity along the trajectory. A smaller $S(v)$ means straighter trajectories. This approach ensures that the difficulty of each Euler step remains consistent, allowing the model to take more steps in more challenging regions. Using the same number of sampling steps, this method outperforms the equal interval approach. The implementation details is provided in Appendix Section A.5. We refer to this approach as equal straightness.
4 Experiments and Analysis
4.1 Experimental Setup
@@
We evaluate RFWave using both Mel-spectrograms and discrete EnCodec tokens as inputs. For Mel-spectrogram inputs, we first benchmark RFWave against existing diffusion vocoders to demonstrate its superiority. We then compare it with widely-used GAN models to highlight its practical applicability and advantages. For discrete EnCodec tokens, we evaluate RFWave’s efficiency in reconstructing high-quality audio from compressed representations across diverse domains. Finally, we conduct ablation studies and further analysis to examine the effects of the individual components of RFWave.
@@
For Mel-spectrogram inputs, we conduct two evaluations. When benchmarking against diffusion vocoders, we train and test on LibriTTS (Zen et al., 2019) (speech), MTG-Jamendo (Bogdanov et al., 2019) (music), and Opencpop (Wang et al., 2022) (vocal) datasets to ensure comprehensive comparison across various audio categories. When comparing to widely used GAN-based models, we train on LibriTTS and test on both LibriTTS and MUSDB18 (Rafii et al., 2017) test subset, evaluating in-domain performance and out-of-domain generalization ability respectively.
For discrete EnCodec token inputs, we follow convention by training a universal model on a large-scale dataset. This dataset combines Common Voice 7.0 (Ardila et al., 2019) and clean data from DNS Challenge 4 (Dubey et al., 2022) for speech, MTG-Jamendo (Bogdanov et al., 2019) for music, and FSD50K (Fonseca et al., 2021) and AudioSet (Gemmeke et al., 2017) for environmental sounds. Recognizing the lack of a comprehensive test set for universal audio codec models, we constructed a unified evaluation dataset comprising 900 test audio samples from 15 external datasets, covering speech, vocals, and sound effects. Detailed information about this test set is provided in Table A.4.
Baseline and Evaluation Metrics
We conducted a comprehensive benchmark comparison of RFWave against a series of state-of-the-art models. For Mel-spectrogram inputs, we utilized PriorGrad (gil Lee et al., 2022) and FreGrad (Nguyen et al., 2024) as baselines for diffusion-based models. Vocos (Siuzdak, 2023) and BigVGAN (gil Lee et al., 2023) served as baselines for GAN-based methods, reflecting their widespread use in real-world applications. For discrete EnCodec token inputs, we compared with EnCodec (Défossez et al., 2022) and MBD (Roman et al., 2023).
We trained PriorGrad and FreGrad on LibriTTS, Opencpop, and MTG-Jamendo using their open-source details. For Vocos, BigVGAN, EnCodec, and MBD, we used public pre-trained models. We followed the authors’ recommended sampling steps: 6 for PriorGrad, 50 for FreGrad, and 20 for MBD. Model sources are in Table A.2.
For objective evaluation, we use ViSQOL (Chinen et al., 2020) to assess perceptual quality, employing speech-mode for 22.05/24 kHz waveforms and audio-mode for 44.1 kHz waveforms. Additional metrics include the Perceptual Evaluation of Speech Quality (PESQ) (Rix et al., 2001), the F1 score for voiced/unvoiced classification (V/UV F1), and the periodicity error (Periodicity) (Morrison et al., 2022). For subjective evaluation, we conduct crowd-sourced assessments, employing a 5-point Mean Opinion Score (MOS) to determine the audio naturalness, ranging from 1 (‘poor/unnatural’) to 5 (‘excellent/natural’).
@@
We use the time-domain model with three enhanced loss functions as the default. The RFWave backbone contains 8 ConvNeXtV2 blocks, and the complex spectrogram is divided into 8 equally spanned subbands. For evaluation, we use 10 sampling steps. Table A.3 lists the parameters used for extracting Mel-spectrograms and complex spectrograms from datasets with different sample rates. Further implementation details and computational resource requirements are provided in the Appendix Sections A.1 and A.2, respectively.
4.2 Experimental Results on Mel-spectrograms Input
Comparasion with Diffusion-based Method
In this portion, we compare RFWave [^3] This evaluation used an earlier RFWave version without the optimized sampling method (equal straightness) from Section 3.3. Despite this, the results convincingly demonstrate RFWave’s superiority over other diffusion vocoders. with PriorGrad and FreGrad on LibriTTS, MTG-Jamendo, and Opencpop datasets. Table 1 presents overall results, while Table 6 provides detailed metrics for each audio category. RFWave consistently outperforms PriorGrad and FreGrad in both objective and subjective metrics. The superior MOS score achieved by RFWave is likely due to its ability to produce clearer and more consistent harmonics, particularly in high-frequency ranges, which contributes to better overall audio quality. A collection of spectrograms and their corresponding analyses can be found in Appendix Section A.6.
Comparasion with GAN-based Method
When comparing RFWave to widely used state-of-the-art GAN-based models (BigVGAN, Vocos), we used models trained on LibriTTS. Besides evaluating on the LibriTTS testset, to assess the models’ robustness and extrapolation capabilities, following the methodology in (gil Lee et al., 2023), we also evaluated these models’ performance on the MUSDB18 dataset. It can be observed in Tables 2 that on the in-domain test set (LibriTTS), RFWave is generally on par with state-of-the-art GAN-based models. However, on the out-of-domain test set (MUSDB18), according to the results in Table 3, RFWave shows significant advantages over BigVGAN and Vocos, even though BigVGAN employed Snake activation (Ziyin et al., 2020) to enhance its out-of-domain data generation capabilities. This demonstrates that RFWave, as a diffusion-type model, offers clear advantages in generalization and robustness compared to GAN-based models. For detail, we observe that GAN-based methods tend to generate horizontal lines in the high-frequency regions of the spectrograms, as exemplified in Figure A.7. In contrast, RFWave consistently produces clear high-frequency harmonics, even when applied to out-of-domain data.
@@ 4.2 Experimental Results on Mel-spectrograms Input
Table 1: Average Mean Opinion Score (MOS) and objective evaluation metrics for RFWave, PriorGrad and FreGrad across various test sets. MOS is provided with 95% confidence interval.
@@
Table 2: MOS and objective evaluation metrics for RFWave, BigVGAN and Vocos on LibriTTS.
Model MOS ↑ PESQ ↑ ViSQOL ↑ V/UV F1 ↑ Periodicity ↓
RFWave 3.82$\pm$0.12 4.304 4.579 0.967 0.091
BigVGAN 3.78$\pm$0.11 4.240 4.712 0.978 0.067
Vocos 3.74$\pm$0.10 3.660 4.696 0.958 0.104
Ground truth 3.91$\pm$0.10 - - - -
Table 3: MOS for RFWave, BigVGAN and Vocos on MUSDB18.
Model Vocals Drums Bass Others Mixture Average
RFWave 3.46$\pm$0.14 3.65$\pm$0.10 3.62$\pm$0.10 3.54$\pm$0.11 4.04$\pm$0.11 3.67$\pm$0.05
BigVGAN 3.42$\pm$0.13 3.68$\pm$0.09 3.50$\pm$0.12 3.33$\pm$0.13 3.58$\pm$0.10 3.51$\pm$0.05
Vocos 3.04$\pm$0.15 3.54$\pm$0.11 2.96$\pm$0.12 2.88$\pm$0.15 3.08$\pm$0.14 3.10$\pm$0.06
Ground truth 3.62$\pm$0.14 3.71$\pm$0.14 3.75$\pm$0.12 3.79$\pm$0.12 4.17$\pm$0.10 3.80$\pm$0.05
4.3 Experimental Results on Discrete Encodec Tokens Input
We evaluated RFWave against EnCodec and MBD across various bandwidths for discrete EnCodec tokens inputs. For RFWave, we used Classifier-Free Guidance (CFG) (Ho & Salimans, 2021) with a 2.0 coefficient. CFG showed little improvement for Mel-spectrogram inputs, but effective for EnCodec tokens, likely due to CFG tending to be more impactful when the input is more compressed.
Table 4 displays the average MOS and objective metrics for EnCodec tokens with varying bandwidths across different auditory contexts. Meanwhile, Table A.6 offers the detailed results for each category. RFWave excels in all metrics, achieving optimal scores, except for ViSQOL. While using a larger bandwidth generally improves performance, for RFWave, an increase in bandwidth from 6.0 kbps to 12.0 kbps results in only a slight enhancement in MOS. EnCodec’s GAN-based decoder attains the highest ViSQOL. Different model families may produce distinct waveform footprints, with ViSQOL showing a subtle bias for GAN-based models.
@@ 4.4 Analysis
We evaluate the model’s performance in both frequency and time domains, then incrementally add the three loss functions and equal straightness to the better-performing model for further analysis. We conduct ablation studies on LJSpeech (Ito & Johnson, 2017), using 250 test sentences for Mel-spectrogram-based waveform reconstruction, summarized in Table 5. Despite the metrics not consistently correlating with human evaluations, they accurately captured the quality improvements brought about by design modifications (Roman et al., 2023). The model operating in the time domain outperforms its frequency domain counterpart, as the ISTFT operation in the former (Figure 1) introduces periodic signals. The inverse Discrete Fourier Transform (IDFT) matrix comprises sinusoidal functions of frequencies $[0,1/N,...,(N-1)/N]$, where $N$ is the number of FFT points (Oppenheim & Schafer, 1975). This mechanism is similar to the Snake activation (Ziyin et al., 2020) in BigVGAN (gil Lee et al., 2023), which also introduces periodic signals into the generator. Energy-balanced loss and equal straightness improves the overall performance. The STFT loss enhances PESQ scores but adversely affects ViSQOL and periodicity error. This occurs because the model tends to prioritize magnitude over phase when applying STFT loss. This trade-off is essential for effectively eliminating artifacts, especially in the presence of background noise. Although overlap loss slightly decreases PESQ, omitting it results in a noticeable transition between subbands. Example spectrograms illustrating the effects of omitting STFT loss and overlap loss are provided in Figures A.8 and A.9, respectively.
Table 5: Objective metrics assess the model’s performance in the frequency domain and the time domain, where three loss functions and equal straightness are applied incrementally.
@@
frequency 3.872 4.430 0.948 0.144
time 4.127 4.551 0.957 0.124
      + energy-balanced loss 4.181 4.598 0.965 0.110
      + overlap loss 4.158 4.599 0.959 0.122
      + STFT loss 4.211 4.578 0.961 0.119
      + equal straightness 4.275 4.654 0.966 0.100
Futher Analysis
Using the model with an 8-layer ConvNeXtV2 backbone, time domain, and three loss functions as the baseline, we first experimented with the DDPM (Ho et al., 2020) approach. Utilizing the noise schedule from DiffWave and PriorGrad, the performance with 50 sampling steps fell short of the baseline, as demonstrated in Table 6. This highlights that Rectified Flow is critical for our model to efficiently generate high-quality audio samples. Next, we experimented with a ResNet (He et al., 2016) backbone with a similar number of parameters. This configuration slightly reduced efficiency and performance in objective metrics (Table 6), indicating that the ConvNeXtV2 backbone enhances both efficiency and audio quality. Finally, we experimented with models of different sizes. Specifically, we tested ConvNextV2 backbones of half and double sizes, with detailed configurations provided in A.1. Increasing the backbone size improves performance but decreases efficiency (Table 6).
Table 6: Objective metrics for further analysis. xRT stands for the speed at which the model can generate speech in comparison to real-time.
Setting GPU xRT ↑ PESQ ↑ ViSQOL ↑ V/UV F1 ↑ Periodicity ↓
baseline 162.59 4.211 4.578 0.961 0.119
DDPM, 50 steps 34.40 2.790 4.499 0.939 0.174
ResNet 150.95 3.865 4.509 0.955 0.128
half size 228.70 4.128 4.553 0.959 0.120
double size 111.62 4.285 4.636 0.968 0.100
4.5 Inference Speed
We perform inference speed benchmark tests using an NVIDIA GeForce RTX 4090 GPU. The implementation was done in PyTorch (Paszke et al., 2019), and no specific hardware optimizations were applied. The inference was carried out with a batch size of 1 sample, utilizing the LJSpeech test set, resampled to model’s sampling rate. Table 7 displays the model size, synthesis speed and GPU memory consumption of the models. RFWave is more than twice as fast as BigVGAN and consumes less GPU memory, thereby eliminating latency as a barrier for practical applications. This speed advantage becomes even more pronounced when synthesizing high-resolution audio (44.1/48 kHz). Frame-level models like RFWave significantly outperform sample-point-based models in high-resolution audio synthesis in terms of speed. RFWave can easily adjust the complex spectrogram’s window lenght and hop length for 44.1 kHz sampling, maintaining low computational complexity, whereas BigVGAN requires additional upsampling blocks. Vocos serves as a strong baseline, given its efficiency in requiring only a single forward pass and operating at the frame level.